All posts by Kona Orlandi

140821 UCSC

Kona got to the roof od ISB at 9:30. No sample had been collected by the collector nor the new funnel apparatus. She washed the collector in place with soap and water, ran a successful test run, then activated the system for fog collection.

140818 LML

Kona arrived to LML at 9 am. The back door had not opened and no sample was collected, a system test proved everything was working fine. The strands appeared clean so she rinsed the apparatus with DI water.

140815 UCSC

Kona arrived at 9 am to a full bottle of sample collected.  The strand frames and top plate were removed for cleaning and everything washed with 409 and DI. The apparatus was reassembled, a blank taken, then set for fog collection.

140814 UCSC

Kona arrived at 9:30 am to find 90 mL of sample collected.  Some dust had accumulated at the bottom of the jar.  The top plate and strand frames were removed and washed at the sink, the rest of the collector was washed in place.  A blank was taken and the system is ready for fog collection.

140725/26/27 LML

On the 25th Kona found that the doors were open but the system had crashed and one of the wires to the small fan was broken.  She reset the system and dismantled the apparatus to wash with 409 at the sink. During the cleaning process the system crashed again.  A system test ran smoothly and the system was reset.

The next morning the system had crashed again after the doors opened.  Kona tested the system manually by increasing the humidity on the RH sensor.  The small fan could not run due to a faulty wire, the doors opened, then there was a click sound but the big fan did not turn on and the system froze.  Kona rinsed the apparatus with DI.  The system was reset again in hope of something different the next day.

The morning of the 27th Kona arrived to find the same situation had occurred.  The apparatus appeared clean so she closed the doors and shut down the system.

140728 UCSC

Kona arrived at 9 am, there was less that 10 mL of water collected so it was discarded.  The apparatus was washed in place with 409 and a thorough DI rinse.  The collecting bottle was rinsed at the sink.  System test was fine, ready for fog collection.

140723 LML

Kona arrived at 11 am.  The doors were open, no sample, the system had crashed and the red wire to the small fan was broken.  The strand frames were taken to the sink and washed with 409 and DI water, the bottle was rinsed with DI water, and the rest of the apparatus was washed in place with 409 and DI water.  During cleaning the system failed again, it was reset once more and activated.

140721 UCSC

Kona arrived at 8 am.  The collector doors were open, no sample collected and the system had crashed.  DI water wash with everything in place; the bottle was rinsed at the sink.   During the wash the system failed again; Kona unplugged the main power, made sure all other connections were secure, then reset and activated the system.  Ready for collection.